# A Diffused Background from Axion-like Particles in the Microwave Sky

### v1

@@ Appendix D Constraints on ALP coupling constant using template matching of foregrounds
Figure 17: Constraints from CMB-S4 using Template Matching at 145 GHz for $\mathrm{z_{min}=0.5}$ and $\mathrm{z_{min}=1}$
Template matching assumes the scaling of foregrounds power spectrums with frequencies and can be used to obtain stronger bounds on the modified coupling constant using a lower beam size. We perform template matching to look for the improvement in constraints.
Figure 18: Constraints from SO using Template Matching at 145 GHz for $\mathrm{z_{min}=0.5}$ and $\mathrm{z_{min}=1}$
Even after masking, the effect of these foregrounds remains even at high latitudes. Thus, a modelling of these foregrounds is required to account for their impact on the power spectrum. These are modelled using high frequencies ($\nu>200$ GHz) for modelling of dust, while low frequencies ($\nu<70$ GHz) for synchrotron emission.
Figure 19: Constraints from CMB-HD using Template Matching at 150 GHz for $\mathrm{z_{min}=0.5}$ and $\mathrm{z_{min}=1}$
We perform a template matching of foregrounds for CMB-S4 configuration to look for any improvement on the bounds on modified coupling constant. We assume we know the scaling of galactic synchrotron and dust emissions with frequencies in the microwave and radio regions of the EM spectra. To model the "s-3" synchrotron model used in our mock data, we use the following equation to account for its curved index [32]:
$\mathrm{C_{\ell}^{syn}(\nu)=A_{\ell}\left(\frac{\nu}{\nu_{0}}\right)^{2\beta_{s}+2Cln(\nu/\nu_{c})}}.$ (D.1)
The curved spectral index shows a steepening or flattening with frequency above the frequency $\mathrm{\nu_{c}}$. The fiducial values used are C = -0.052, $\mathrm{\nu_{c}=23}$ GHz and $\mathrm{\beta_{s}}$ = -3.
The modified black-body function with $\mathrm{\beta_{d}=1.58}$ is used to fit the dust model "d-3":
$\mathrm{(C_{\ell})_{dust}=A_{\ell}\nu^{2\beta_{d}}B^{2}_{\nu}(T).}$ (D.2)
By modelling the foreground emissions at low (for synchrotron) and high (for dust) frequencies, the scaling with frequency can be used to obtain the contribution of synchrotron and dust at the frequencies 90 - 160 GHz. This method assumes the scaling of foregrounds with frequencies for multipole range $\ell>\mathrm{\ell_{max}}$ corresponding to the beam of the instrument at the matched frequency of 145 GHz for CMB-S4 and SO, while 150 GHz for CMB-HD.
@@
The constraints obtained on the modified ALP coupling constant A are shown for various CMB surveys for minimum redshifts $\mathrm{z_{min}=0.5}$ and $\mathrm{z_{min}=1}$ in Figs. 17 (CMB-S4: A < 0.654 with $\mathrm{z_{min}=0.5}$ and A < 0.734 with $\mathrm{z_{min}=1}$ ), 18 (SO: A < 2.005 with $\mathrm{z_{min}=0.5}$ and A < 2.325 with $\mathrm{z_{min}=1}$ ) and 19 (CMB-HD: A < 0.115 with $\mathrm{z_{min}=0.5}$ and A < 0.131 with $\mathrm{z_{min}=1}$ ). The constraints are expected with CMB-HD giving the tightest constraints. The constraints from template matching are stronger, as compared to ILC for all the three detectors (SO, CMB-S4 and CMB-HD).
The converted constraints (95% C.I.) on ALP coupling constant $\mathrm{g_{a\gamma}}$ using template matching are shown in Fig.20. With template matching, we will be able to get the following bounds with $\mathrm{z_{min}=0.5}$: SO - $\mathrm{g_{a\gamma}<1.416\times 10^{-12}\,GeV^{-1}}$ ; CMB-S4 - $\mathrm{g_{a\gamma}<8.087\times 10^{-13}\,GeV^{-1}}$; CMB-HD - $\mathrm{g_{a\gamma}<3.391\times 10^{-13}\,GeV^{-1}}$, while the bounds with $\mathrm{z_{min}=1}$ go as: SO - $\mathrm{g_{a\gamma}}<1.525\times 10^{-12}$ GeV-1 ; CMB-S4 - $\mathrm{g_{a\gamma}<8.567\times 10^{-13}\,GeV^{-1}}$; CMB-HD - $\mathrm{g_{a\gamma}<3.619\times 10^{-13}\,GeV^{-1}}$.
Figure 20: Converted ALP coupling constant constraints for various detectors using Template Matching for $\mathrm{z_{min}=0.5}$ and $\mathrm{z_{min}=1}$
Both template matching and ILC assume the shape of the foregrounds power spectrum for $\mathrm{\ell>\ell_{max}}$, but since the beam size is larger for ILC (lower $\mathrm{\ell_{max}}$), template matching provides better bounds with a higher $\mathrm{\ell_{max}}$ corresponding to a lower beam size at the matched frequency for all detectors, but is less reliable as the modelling of foregrounds at different frequencies may not be precise leading to bias in the constraints.

### v3

@@ Appendix D Constraints on ALP coupling constant using template matching of foregrounds
Figure 17: Constraints from CMB-S4 using Template Matching at 145 GHz for $z_{\mathrm{min}}=0.5$ and $z_{\mathrm{min}}=1$
Template matching assumes the scaling of the foregrounds power spectra with frequencies and can be used to obtain stronger bounds on the modified coupling constant using a lower beam size. We perform template matching to look for the improvement in constraints.
Figure 18: Constraints from SO using Template Matching at 145 GHz for $z_{\mathrm{min}}=0.5$ and $z_{\mathrm{min}}=1$
Even after masking, the effect of these foregrounds remains even at high latitudes. Thus, a modelling of these foregrounds is required to account for their impact on the power spectrum. These are modelled using high frequencies ($\nu>200$ GHz) for dust, while low frequencies ($\nu<70$ GHz) for synchrotron emission.
Figure 19: Constraints from CMB-HD using Template Matching at 150 GHz for $z_{\mathrm{min}}=0.5$ and $z_{\mathrm{min}}=1$
We perform a template matching of foregrounds for CMB-S4 configuration to look for any improvement in the bounds on modified coupling constant. We assume we know the scaling of galactic synchrotron and dust emissions with frequencies in the microwave and radio regions of the EM spectra. To model the "s-3" synchrotron model used in our mock data, we use the following equation to account for its curved index [55]:
$C_{\ell}^{\mathrm{syn}}(\nu)=A_{\ell}\left(\frac{\nu}{\nu_{0}}\right)^{2\beta_{s}+2C\ln(\nu/\nu_{c})}.$ (D.1)
The curved spectral index shows a steepening or flattening with frequency above the frequency $\nu_{c}$. The fiducial values used are $C=-0.052$, $\nu_{c}=23$ GHz and $\beta_{s}=-3$.
The modified black-body function with $\beta_{d}=1.58$ is used to fit the dust model "d-3":
$C_{\ell}^{\mathrm{dust}}(\nu)=A_{\ell}\nu^{2\beta_{d}}B^{2}_{\nu}(T).$ (D.2)
By modelling the foreground emissions at low (for synchrotron) and high (for dust) frequencies, the scaling with frequency can be used to obtain the contribution of synchrotron and dust at the frequencies 90 - 160 GHz. This method assumes the scaling of foregrounds with frequencies for multipole range $\ell>\ell_{\mathrm{max}}$ corresponding to the beam of the instrument at the matched frequency of 145 GHz for CMB-S4 and SO, while 150 GHz for CMB-HD.
@@
The constraints obtained on the modified ALP coupling constant $A$ are shown for various CMB surveys for minimum redshifts $z_{\mathrm{min}}=0.5$ and $z_{\mathrm{min}}=1$ in Figs. 17 (CMB-S4: $A<0.654$ with $z_{\mathrm{min}}=0.5$ and $A<0.734$ with $z_{\mathrm{min}}=1$ ), 18 (SO: $A<2.005$ with $z_{\mathrm{min}}=0.5$ and $A<2.325$ with $z_{\mathrm{min}}=1$ ) and 19 (CMB-HD: $A<0.115$ with $z_{\mathrm{min}}=0.5$ and $A<0.131$ with $z_{\mathrm{min}}=1$ ). The constraints are expected with CMB-HD giving the tightest constraints. The constraints from template matching are stronger, as compared to ILC for all three detectors (SO, CMB-S4, and CMB-HD).
The converted constraints (95% C.I.) on ALP coupling constant $g_{a\gamma}$ using template matching are shown in Fig.20. With template matching, we will be able to get the following bounds with $z_{\mathrm{min}}=0.5$: SO - $g_{a\gamma}<1.416\times 10^{-12}\,\mathrm{GeV^{-1}}$; CMB-S4 - $g_{a\gamma}<8.087\times 10^{-13}\,\mathrm{GeV^{-1}}$; CMB-HD - $g_{a\gamma}<3.391\times 10^{-13}\,\mathrm{GeV^{-1}}$, while the bounds with $z_{\mathrm{min}}=1$ go as:
$\mathrm{SO}:g_{a\gamma}<1.525\times 10^{-12}\mathrm{GeV^{-1}};$
$\mathrm{CMB-S4}:g_{a\gamma}<8.567\times 10^{-13}\,\mathrm{GeV^{-1}};$
$\mathrm{CMB-HD}:g_{a\gamma}<3.619\times 10^{-13}\,\mathrm{GeV^{-1}}.$
Figure 20: Converted ALP coupling constant constraints for various detectors using Template Matching for $z_{\mathrm{min}}=0.5$ and $z_{\mathrm{min}}=1$
Both template matching and ILC assume the shape of the foregrounds power spectrum for $\ell>\ell_{\mathrm{max}}$, but since the beam size is larger for ILC (lower $\ell_{\mathrm{max}}$), template matching provides better bounds with a higher $\ell_{\mathrm{max}}$ corresponding to a lower beam size at the matched frequency for all detectors, but is less reliable as the modelling of foregrounds at different frequencies may not be precise leading to bias in the constraints.